# Interface instability of two-phase flow in a three-dimensional porous medium

### v1

@@ I Introduction
Fluid invasion describes the process by which a pressure gradient impels one fluid to displace another within a porous medium. The stability and evolution of the resulting fluid interface depends on the balance of gravitational, viscous, and capillary forces, as well as wetting properties and any changes in the solid structure of the porous medium [34, 23, 24, 26, 25, 27, 28, 29]. In the early 2D experiments on fluid invasion [35, 25], glass beads were sandwiched between glass plates in a porous version of the classic Hele-Shaw cell [36], and the resulting fluid dynamics were imaged. These experiments produced fluid interfaces whose geometries varied with the applied pressure gradient from compact and non-fractal to ramified and fractal [25, 37, 35, 26, 38]. Depending on the balance of the controlling forces, the immiscible-fluid interfaces or “fronts” can be either stable, with a relatively compact invasion structure and a width that stabilizes at a constant value, or unstable, with a width that grows continually and a fractal morphology, appearing as a collection of overlaid fingers [39, 30, 25, 26, 40]. The dimensionless fluctuation number $F$ was introduced in [39] within a modified percolation theory of front geometry, and it weighs the relative importance of viscous and gravitational forces against the capillary pressure-threshold fluctuations to characterize front width and stability [30]. Assuming that the width of an invasion front scales with the typical size of defending-fluid clusters, one predicts a relation between the front width and the fluctuation number, with an unstable interface for $F<0$ and a stable one for $F>0$ [25, 26, 40, 30, 41]. Although this fluctuation-number theory nicely describes invasion-front stability and width in 2D experiments, verifying whether the theory is easily adapted to describe invasion fronts from 3D experiments remains an important task [30].
@@
In the present work, we apply RIM-LIF imaging to study the geometry and stability of an initially planar front progressing through a beadpack. Experiments in which the front width reaches a quasi-constant value in time are considered stable, while those in which the width grows continually over the resolved length- and time-scales are considered unstable. While our earlier work evaluates a point-source injection and thereby samples a wide range of front velocities as the front expands and slows down [31, 32, 33], this work evaluates a planar injection, selected to investigate the front width evolution as a function of global flow rate and to describe an expected transition from stable to unstable invasion. We evaluate the stability of the immiscible fluid interfaces as they move during the invasion using time-resolved three-dimensional images of the dynamics. Specifically, we consider a fluid with higher density and viscosity invading one with lower density and viscosity. Depending on the flow rate, viscosity either stabilizes the front or gravity destabilizes it, analogous to the Rayleigh-Taylor instability [51]. The image data are complemented by time series of the fluid pressures at the inlet and outlet of the experimental cell, giving a comprehensive picture of interface stability during 3D invasion flows.
@@ II.2 Porous-medium preparation and fluid characteristics
In our two-phase flow experiments, the interfacial tension between the two fluids and the wetting characteristics of the two fluids on the solid material controls the capillary pressures required to invade the individual pores, so these factors could affect the invasion front morphology. With a tensiometer we measured the fluids’ interfacial tension to be $\gamma=(16.4\pm 0.4)$ mN/m, and found neutral fluid-solid wetting angles for both experimental fluids. In trial experiments, using glycerol as the invading phase showed that it appeared to be the non-wetting fluid, passing through pore throats with a convex profile, using oil as the invading phase showed that oil also appeared to be the non-wetting phase, with a similar convex profile in the pore throats. We therefore consider the wetting conditions to be of secondary importance to the invasion front dynamics in what follows, at least when compared to the balance of viscous and gravitational forces at the front.
@@ III Theoretical criterion for front instability
Now we consider a two-phase configuration in which one fluid invades a volume initially saturated with another. We consider an initially flat front at $z=z_{0}$ and perturb one region of it by a small distance $a$, comparable to the size of an individual pore. Evaluating the minimal flow rate for which this perturbation grows provides an estimate of $Q_{c}$. Using Eq. (4), the pressure changes in the invading fluid (1) and the defending fluid (2) due to the perturbation can be expressed as
$\begin{split}\Delta p_{1}&=p_{1}(z_{0}+a)-p_{1}(z_{0})=\left(\rho_{1}g-\frac{Q\mu_{1}}{A\kappa_{TP}(Q)}\right)a\;,\\ \Delta p_{2}&=p_{2}(z_{0}+a)-p_{2}(z_{0})=\left(\rho_{2}g-\frac{Q\mu_{2}}{A\kappa_{SP}}\right)a\;,\end{split}$ (5)
where $\kappa_{SP}$ and $\kappa_{TP}(Q)$ are the respective single- and two-phase permeabilities [33]. Here, we allow for $\kappa_{TP}(Q)$ to explicitly depend on the flow rate $Q$, reflecting an expectation that clusters of defending fluid trapped behind the main invading front will effectively reduce the permeability for the invading phase, meaning that, in general, $\kappa_{TP}(Q)<\kappa_{SP}$ [56, 57, 58].
The capillary pressure across the interface between fluids 1 and 2 is $p^{c}=p_{1}-p_{2}$, so the capillary pressure change induced by the perturbation at a given flow rate $Q$ is $\Delta p^{c}(Q)=\Delta p_{1}-\Delta p_{2}$, or using Eq. (5),
$\Delta p^{c}(Q)=\left[(\rho_{1}-\rho_{2})g-\frac{Q}{A}\left(\frac{\mu_{1}}{\kappa_{TP}(Q)}-\frac{\mu_{2}}{\kappa_{SP}}\right)\right]a\;.$ (6)
Because the invading fluid (1) has higher viscosity and density than the defending fluid (2), $\Delta p^{c}$ is increased by the gravitational term and decreased by the viscous term. The perturbation is unstable and tends to grow when gravity dominates, $\Delta p^{c}(Q)>0$, and conversely it is stable and tends to shrink when viscosity dominates, $\Delta p^{c}(Q)<0$. The rate $Q_{c}$, below which fronts destabilize, therefore follows from $\Delta p_{c}(Q_{c})=0$. Noting that $\mu_{1}\gg\mu_{2}$ (see Table 1), we obtain the condition
@@
for the critical flow rate $Q_{c}$. Fronts are expected to be unstable for all $Q\lesssim Q_{c}$, where $Q_{c}$ is the solution to Eq. (7). In Sec. IV.4, we study the dependence of $\kappa_{TP}(Q)$ on $Q$ to predict $Q_{c}$ from Eq. (7), and we compare the prediction with the observed front-stability trends.
@@ IV Results
Table 2: Summary of experimental results. The Stability column indicates whether the front was classified as stable or unstable for each experiment. Experiments were considered unstable when the best-fit slope of $W/a$ versus the injected volume exceeded $0.005$ ml-1, see Sec. IV.1. Our classification is largely insensitive to this threshold, although changes could modify the assessed stability of some experiments. The Crystal fraction represents the proportion of beads participating in crystallites in each experiment, as described in Sec. IV.5. $\frac{dp_{b}-dp_{t}}{dt}$ is the averaged time derivative of the pressure difference across the cell, as described in Sec. IV.3, with errors estimated as the maximum possible rate-of-change of hydrostatic pressure in the outlet basin ($\rho_{2}gQ/A_{b}$, with $A_{b}$ the basin area). Pressure measurements were not performed for Exps. C and G, denoted by hyphens.
ID $Q$ [ml/min] Stability Crystal fraction $\frac{dp_{b}-dp_{t}}{dt}$ [Pa/s]
@@ IV.1 Front instability: contributions of fingering and pinning
The experimental images indicate two different modes by which fronts destabilize, as shown in Fig. 5. In the first mode, which we refer to as fingering, thin regions of invading fluid increasingly protrude ahead of the main front, contributing forward skew to the front width distribution. In the second mode, referred to as pinning, regions of invading fluid lag behind the main front, contributing backward skew in the front distribution. When the front distribution is symmetrical, neither fingering nor pinning dominates, and we refer to the front distribution as neutral. Snapshots of representative neutral, fingered, and pinned fronts are displayed in Figs. 5(a-c).
To quantify the relative influence of fingering and pinning on the observed front dynamics, we evaluate contributions to front-width growth by the positive and negative front widths, $W_{+}$ and $W_{-}$, defined in Eqs. (2) and (3) and plotted in Fig. 4(b) and (c). The differing contributions of fingering and pinning to skew in the front width distributions are displayed in Fig. 6, where a stable front (Exp. C) is compared to fronts destabilized predominantly by fingering (Exp. H) and pinning (Exp. Q), respectively. In the figure, the front distribution is scaled to show the proportion of front voxels at a given vertical position.
@@ IV.3 Analysis of pressure readings
$\frac{d(p_{b}-p_{t})}{dt}\approx v_{f}(Q)\frac{d(p_{b}-p_{t})}{dz}=v_{f}(Q)\left[(\rho_{1}-\rho_{2})g-\frac{Q}{A}\left(\frac{\mu_{1}}{\kappa_{TP}(Q)}-\frac{\mu_{2}}{\kappa_{SP}}\right)\right].$ (11)
This result shows the same balance of viscous and gravitational pressures encoded in Eq. (6). We therefore expect a stable front when $\frac{d(p_{b}-p_{t})}{dt}<0$ and an unstable front otherwise. When the front is unstable, Eqs. (10) and (11) are no longer valid, since the front can no longer be represented as a sharp interface.
We measured pressure signals in all experiments except Exps. C and G. Figure 9 shows $p_{t}$ and $p_{b}$ from three representative experiments. In each experiment, the pressures follow a three-stage response, corresponding to (I) experiment initialization, in which the area above the mesh screen fills with invading fluid; (II) initial invasion, in which the fluid permeates the cell at the maximum injection rate into the imaging region; and (III) interface dynamics, when the pressures evolve under the imposed flow rate, $Q$. If the interface is stable, the final stage of pressure evolution (III) is approximated by Eqs. (10) and (11). According to Eq. (11), the inequality $\frac{dp_{b}-dp_{t}}{dt}<0$ provides a pressure-based criterion for front stability. By interpreting the time derivative of the pressure difference in Table 2, we could use this inequality to qualitatively estimate the critical flow rate $Q_{c}$. Instead, we evaluate $Q_{c}$ with a quantitative method in the following section.
@@ IV.4 Flow-dependent two-phase permeability and the critical flow rate
Figure 10: Panel (a) shows the mobility $\kappa_{TP}(Q)/\mu_{1}$ for different flow rates $Q$. These data are calculated from the experimental data and Eq. (11), as described in Sec. IV.4. Each experiment provides one data point, and the points have colors representing whether the experiment was classified as stable (green), unstable (blue), or as an outlier as far as the fitting is concerned (gray), as described in the text. A linear fit $\kappa_{TP}/\mu_{1}=B_{1}+B_{2}Q$ is shown as a solid black line, while the power-law fit $\kappa_{TP}/\mu_{1}=B_{1}^{\prime}+B_{2}^{\prime}Q^{0.5}$ is shown for comparison as a dotted gray line. The best-fit parameters are provided in the text. Panel (b) shows the averaged time derivative of $p_{b}-p_{t}$ versus the flow rate, with Eq. (11) evaluated for both linear and power-law fits of $\kappa_{TP}$, providing nearly identical results. The solid black line indicates the solution $Q_{c}=(5.5\pm 2.9)$ ml/min of Eq. (6) using the best-fit parameters with the linear model, while the shaded region is the 95% confidence interval. This solution is equivalent to $d(p_{b}-p_{t})/dt=0$. Two experiments were excluded from the fitting (Exps. G and Q) based on irregularities in their porous geometries, discussed in Sec. IV.5.
@@
A power-law dependence of $\kappa_{TP}$ on $Q$ with an exponent $\alpha=0.5$ has been measured in steady-state experiments, where both fluids are injected at the same time [57, 58], although this scenario differs from the experiments described here. Due to the scatter in the experimental data, we are not able to decide the functional dependence of $\kappa_{TP}/\mu_{1}$ on Q. The scatter reflects variations in the porosity and arrangement of grains in the porous structure, giving rise to variations in the permeability and pressure measurements from sample to sample. This produces relative uncertainties in $\frac{d(p_{b}-p_{t})}{dt}$ and $\kappa/\mu_{1}$ of the order of 20%. Nevertheless, both cases $\alpha=1$ and $\alpha=0.5$ fit the experimental data better than choosing a constant value for $\kappa_{TP}(Q)/\mu_{1}$, as seen in Fig. 10(a).
Using the best fit model to $\kappa_{TP}(Q)$ we plot Eq. (11) in Fig. 10(b), showing consistency with the experimental data. We also use this relation to identify the critical flow rate from the pressure data, which complements our earlier estimate from the image data in Sec. IV.1. Using Eq. (11) to evaluate $d(p_{b}-p_{t})/dt=0$, which is equivalent to solving Eq. (7), provides $Q_{c}=(5.5\pm 2.9)$ ml/min, with listed uncertainties from 95% confidence bounds. This pressure-derived value of $Q_{c}$ agrees with our image-derived estimate within the experimental uncertainty.

### v2

@@ I Introduction
Fluid invasion describes the process by which a pressure gradient impels one fluid to displace another within a porous medium. The stability and evolution of the resulting fluid interface depends on the balance of gravitational, viscous, and capillary forces, as well as wetting properties and any changes in the solid structure of the porous medium [34, 23, 25, 27, 24, 26, 28, 35, 29]. Most of the early experiments on fluid invasion imaged fluid displacement patterns in quasi-2D systems, such as etched glass networks [36, 37] and beads sandwiched between glass plates [38, 25] in a porous analogue of the Hele-Shaw cell [39]. Other early studies imaged 2D slices through 3D displacement patterns [40, 41, 42, 27, 43], at a time when computational power and imaging techniques did not allow detailed resolution of the full 3D fluid-invasion bodies. These experiments produced fluid interfaces whose geometries varied with the applied pressure gradient from compact and non-fractal to ramified and fractal [25, 44, 38, 26, 45]. Depending on the balance of the controlling forces, the immiscible-fluid interfaces or “fronts” can be either stable, with a relatively compact invasion structure and a width that stabilizes at a constant value, or unstable, with a width that grows continually and a fractal morphology, appearing as a collection of overlaid fingers [46, 30, 25, 26, 47]. The dimensionless fluctuation number $F$ was introduced in [46] within a modified percolation theory of front geometry, and it weighs the relative importance of viscous and gravitational forces against the capillary pressure-threshold fluctuations to characterize front width and stability [30]. Assuming that the width of an invasion front scales with the typical size of defending-fluid clusters, one predicts a relation between the front width and the fluctuation number, with an unstable interface for $F<0$ and a stable one for $F>0$ [25, 26, 47, 30, 48]. Although this fluctuation-number theory nicely describes invasion-front stability and width in 2D experiments, verifying whether the theory is easily adapted to describe invasion fronts from 3D experiments remains an important task [30].
@@
In the present work, we apply RIM-LIF imaging to study the geometry and stability of an initially planar front progressing through a beadpack. Experiments in which the front width reaches a quasi-constant value in time are considered stable, while those in which the width grows continually over the resolved length- and time-scales are considered unstable. While our earlier work evaluates a point-source injection and thereby samples a wide range of front velocities as the front expands and slows down [31, 32, 33], this work evaluates a planar injection, selected to investigate the front width evolution as a function of global flow rate and to describe an expected transition from stable to unstable invasion. We evaluate the stability of the immiscible fluid interfaces as they move during the invasion using time-resolved three-dimensional images of the dynamics. Specifically, we consider a fluid with higher density and viscosity invading one with lower density and viscosity from the top. Depending on the flow rate, viscosity either stabilizes the front or gravity destabilizes it, analogous to the Rayleigh-Taylor instability [57]. We regard the results in light of numerous studies of the interplay between capillary, viscous, and gravitational forces [34, 58, 42, 25, 27, 43, 46, 26, 28, 17, 47, 59]. The image data are complemented by time series of the fluid pressures at the inlet and outlet of the experimental cell, giving a comprehensive picture of interface stability during 3D invasion flows.
@@ II.2 Porous-medium preparation and fluid characteristics
In our two-phase flow experiments, the interfacial tension between the two fluids and the wetting characteristics of the two fluids on the solid material controls the capillary pressures required to invade the individual pores, so these factors could affect the invasion front morphology. With a tensiometer we measured the fluids’ interfacial tension to be $\gamma=(16.4\pm 0.4)$ mN/m, and found neutral fluid-solid wetting angles for both experimental fluids. In trial experiments, close inspection of the liquid–liquid interface showed that both glycerol-displacing-oil and oil-displacing-glycerol events produced meniscus profiles characteristic of drainage, suggesting that the displacing fluid behaved as the non-wetting phase in both cases. We therefore conclude that neither fluid exhibited a clearly dominant wetting tendency and that the system exhibits near-neutral wettability, with conditions remaining consistent throughout the experiments. Given this neutral wettability condition, we chose not to emphasize wetting effects in the analysis that follows, focusing instead on the balance between viscous and gravitational forces at the invasion front. However, for a comprehensive understanding of invasion morphologies, wettability remains an important factor [41, 61].
@@ III Theoretical criterion for front instability
Now we consider a two-phase configuration in which one fluid invades a volume initially saturated with another. We consider an initially flat front at $z=z_{0}$ and perturb one region of it by a small distance $a$, comparable to the size of an individual pore. Evaluating the minimal flow rate for which this perturbation grows provides an estimate of $Q_{c}$. Using Eq. (4), the pressure changes in the invading fluid (1) due to the perturbation can be expressed as
$\Delta p_{1}=p_{1}(z_{0}+a)-p_{1}(z_{0})=\left(\rho_{1}g-\frac{Q\mu_{1}}{A\kappa_{TP}(Q)}\right)a$ (5)
The corresponding pressure change in the displaced fluid (2) of an unperturbed interface is
$\Delta p_{2}=p_{2}(z_{0}+a)-p_{2}(z_{0})=\left(\rho_{2}g-\frac{Q\mu_{2}}{A\kappa_{SP}}\right)a\;.$ (6)
Here $\kappa_{SP}$ and $\kappa_{TP}(Q)$ are the respective single- and two-phase permeabilities [33]. Here, we allow $\kappa_{TP}(Q)$ to explicitly depend on the flow rate $Q$, reflecting the expectation that clusters of defending fluid trapped behind the main invading front will effectively reduce the permeability for the invading phase, meaning that, in general, $\kappa_{TP}(Q)<\kappa_{SP}$ [65, 66, 67].
The increased pressure acting on the fluid-fluid interface induced by the perturbation at a given flow rate $Q$ is $\Delta p^{d}(Q)=\Delta p_{1}-\Delta p_{2}$, or using Eq. (6),
$\Delta p^{d}(Q)=\left[(\rho_{1}-\rho_{2})g-\frac{Q}{A}\left(\frac{\mu_{1}}{\kappa_{TP}(Q)}-\frac{\mu_{2}}{\kappa_{SP}}\right)\right]a\;.$ (7)
If the net pressure $\Delta p^{d}(Q)$ is positive, any small displacement will be amplified, resulting in instability. In contrast, if $\Delta p^{d}(Q)$ is negative, the perturbation will be dampened. This holds true regardless of the surface tension value.
This result can be rewritten in terms of an effective capillary number, $\mathrm{Ca}_{\text{eff}}$ and Bond number, $\mathrm{Bo}$ as
$\Delta p^{d}(Q)=\frac{\gamma}{a}\left[\mathrm{Bo}-\mathrm{Ca}_{\text{eff}}(Q)\right],$ (8)
where
$\mathrm{Bo}=\frac{a^{2}(\rho_{1}-\rho_{2})g}{\gamma}\quad\text{ and }\quad\mathrm{Ca}_{\text{eff}}(Q)=\frac{a^{2}}{\gamma}\cdot\frac{Q}{A}\left(\frac{\mu_{1}}{\kappa_{TP}(Q)}-\frac{\mu_{2}}{\kappa_{SP}}\right).$ (9)
In this formulation, we expect the transition from stable to unstable to occur when $\mathrm{Ca}_{\text{eff}}=\mathrm{Bo}$.
In our experiments, because the invading fluid (1) has higher viscosity and density than the defending fluid (2), $\Delta p^{d}$ increases by the gravitational term and decreases by the viscous term. The perturbation is unstable and tends to grow when gravity dominates, $\Delta p^{d}(Q)>0$, and conversely it is stable and tends to shrink when viscosity dominates, $\Delta p^{d}(Q)<0$. The rate $Q_{c}$, below which fronts destabilize, therefore follows from $\Delta p^{d}(Q_{c})=0$. Noting that $\mu_{1}\gg\mu_{2}$ (see Table 1), we obtain the condition
@@
for the critical flow rate $Q_{c}$. Fronts are expected to be unstable for all $Q\lesssim Q_{c}$, where $Q_{c}$ is the solution to Eq. (10). This result aligns with the findings from the linear stability analysis conducted by Saffman and Taylor [34] and Chuoke et al. [68]. In Sec. IV.4, we study the dependence of $\kappa_{TP}(Q)$ on $Q$ to predict $Q_{c}$ from Eq. (10), and we compare the prediction with the observed front-stability trends. In Sec. V.4 we discuss the theory further, and expand on how the capillary threshold distribution affects the front width.
@@ IV Results
Table 2: Summary of experimental results. The Stability column indicates whether the front was classified as stable or unstable for each experiment. Experiments were considered unstable when the best-fit slope of $W/a$ versus the injected volume exceeded $0.005$ ml-1, see Sec. IV.1. Our classification is largely insensitive to this threshold, although changes could modify the assessed stability of some experiments. The “Crystal fraction” represents the proportion of beads participating in crystallites in each experiment, as described in Sec. IV.5. $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t$ is the averaged time derivative of the pressure difference across the cell, as described in Sec. IV.3, with errors estimated as the maximum possible rate-of-change of hydrostatic pressure in the outlet basin ($\rho_{2}gQ/A_{b}$, with $A_{b}$ the basin area). Pressure measurements were not performed for Exps. C and G, denoted by hyphens.
ID $Q$ [ml/min] Stability Crystal fraction $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t$ [Pa/s]
@@ IV.1 Front instability: contributions of fingering and pinning
The experimental images revealed two different mechanisms by which fronts destabilized, which we referred to as fingering and pinning (Sec. IV.1 and Fig. 5). Fingering appears as narrow plumes of invading fluid protruding in the direction of the flow. This instability originates from the interplay between the pressure gradients within the fluid phases due to viscous and gravitational forces, and the capillary-pressure thresholds for interface displacement through individual pores. If a sufficiently large positive capillary pressure gradient forms along the fluid interface in the flow direction, fingering occurs, and this mechanism has long been understood to control front instability in 2D experiments [38, 30, 26, 7, 48]. Conversely, a negative capillary pressure gradient in the direction of flow stabilizes the invasion front and suppresses the development of fingering [58, 25, 69, 47, 30, 59, 48, 70, 71].
We define pinning as the scenario in which the invading front bypasses a correlated region of the pore space characterized by unusually low permeability [see Fig. 5(c)]. This occurs when the interface encounters an area with significantly smaller pores, where the local capillary pressure is typically insufficient to overcome the higher entry thresholds imposed by the tighter geometry. As a result, the front appears to become pinned in that region, while invasion proceeds preferentially through the surrounding wider pore throats. Most of our unstable 3D experiments show some evidence of pinning, which we identify through the relatively large contribution of backward-skewed regions in the front distribution to the overall front width (see Fig. 6). Previous numerical studies of gravity-driven invasion percolation have similarly demonstrated that introducing large-scale structural correlations in the porous medium can substantially alter front progression and the resulting residual fluid saturation after drainage [72]. In our case, pinning is associated with regions of correlated crystalline structures much larger than individual pores that impede the progression of the invasion front due to their higher capillary pressure thresholds. See Sec. IV.5 for further discussion of the correlated crystalline structures.
When the front distribution is symmetrical, neither fingering nor pinning dominates, and we refer to the front distribution as neutral. For clarity, this does not imply that the front is stable, as the front width might still continue growing indefinitely despite having a symmetrical distribution. Snapshots of representative neutral, fingered, and pinned fronts are displayed in Figs. 5(a-c).
To quantify the relative influence of fingering and pinning on the observed front dynamics, we evaluate contributions to front-width growth by the positive and negative front widths, $W_{+}$ and $W_{-}$, defined in Eqs. (2) and (3) and plotted in Fig. 4(b) and (c). The different contributions of fingering and pinning to skew in the front width distributions are displayed in Fig. 6, where a stable front (Exp. C) is compared to fronts predominantly destabilized by fingering (Exp. H) and pinning (Exp. Q), respectively. In the figure, the front distribution is scaled to show the proportion of front voxels at a given vertical position.
@@ IV.3 Analysis of pressure readings
$\frac{\mathrm{d}(p_{b}-p_{t})}{\mathrm{d}t}\approx v_{f}(Q)\frac{\mathrm{d}(p_{b}-p_{t})}{\mathrm{d}z}=v_{f}(Q)\left[(\rho_{1}-\rho_{2})g-\frac{Q}{A}\left(\frac{\mu_{1}}{\kappa_{TP}(Q)}-\frac{\mu_{2}}{\kappa_{SP}}\right)\right].$ (14)
This result shows the same balance of viscous and gravitational pressures encoded in Eq. (7). We therefore expect a stable front when $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t<0$ and an unstable front otherwise. When the front is unstable, Eqs. (13) and (14) are no longer valid, since the front can no longer be represented as a sharp interface.
We measured pressure signals in all experiments except Exps. C and G. Figure 9 shows $p_{t}$ and $p_{b}$ from three representative experiments. In each experiment, the pressures follow a three-stage response, corresponding to (I) experiment initialization, in which the area above the mesh screen fills with invading fluid; (II) initial invasion, in which the fluid permeates the cell at the maximum injection rate into the imaging region; and (III) interface dynamics, when the pressures evolve under the imposed flow rate, $Q$. If the interface is stable, the final stage of pressure evolution (III) is approximated by Eqs. (13) and (14). According to Eq. (14), the inequality $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t<0$ provides a pressure-based criterion for front stability. By interpreting the time derivative of the pressure difference in Table 2, we could use this inequality to qualitatively estimate the critical flow rate $Q_{c}$. Instead, we evaluate $Q_{c}$ with a quantitative method in the following section.
@@ IV.4 Flow-dependent two-phase permeability and the critical flow rate
Figure 10: Panel (a) shows the mobility $\kappa_{TP}(Q)/\mu_{1}$ for different flow rates $Q$. These data are calculated from the experimental data and Eq. (14), as described in Sec. IV.4. Each experiment provides one data point, and the points have colors representing whether the experiment was classified as stable (green), unstable (blue), or as an outlier as far as the fitting is concerned (gray), as described in the text. A linear fit $\kappa_{TP}/\mu_{1}=B_{1}+B_{2}Q$ is shown as a solid black line, while the power-law fit $\kappa_{TP}/\mu_{1}=B_{1}^{\prime}+B_{2}^{\prime}Q^{0.5}$ is shown for comparison as a dotted gray line. The best-fit parameters are provided in the text. Panel (b) shows the averaged time derivative of $p_{b}-p_{t}$ versus the flow rate, with Eq. (14) evaluated for both linear and power-law fits of $\kappa_{TP}$, providing nearly identical results. The solid black line indicates the solution $Q_{c}=(5.5\pm 2.9)$ ml/min of Eq. (7) using the best-fit parameters with the linear model, while the shaded region is the 95% confidence interval. This solution is equivalent to $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t=0$. Two experiments were excluded from the fitting (Exps. G and Q) based on irregularities in their porous geometries, discussed in Sec. IV.5.
@@
A power-law dependence of $\kappa_{TP}$ on $Q$ with an exponent $\alpha=0.5$ has been measured in steady-state experiments, where both fluids are injected at the same time [66, 67], although this scenario differs from the experiments described here. Due to the scatter in the experimental data, we are not able to decide the functional dependence of $\kappa_{TP}/\mu_{1}$ on Q. The scatter reflects variations in the porosity and arrangement of grains in the porous structure, giving rise to variations in the permeability and pressure measurements from sample to sample. This produces relative uncertainties in $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t$ and $\kappa/\mu_{1}$ of the order of 20%. Nevertheless, both cases $\alpha=1$ and $\alpha=0.5$ fit the experimental data better than choosing a constant value for $\kappa_{TP}(Q)/\mu_{1}$, as seen in Fig. 10(a).
Using the best fit model to $\kappa_{TP}(Q)$ we plot Eq. (14) in Fig. 10(b), showing consistency with the experimental data. We also use this relation to identify the critical flow rate from the pressure data, which complements our earlier estimate from the image data in Sec. IV.1. Using Eq. (14) to evaluate $\mathrm{d}(p_{b}-p_{t})/\mathrm{d}t=0$, which is equivalent to solving Eq. (10), provides $Q_{c}=(5.5\pm 2.9)$ ml/min, with listed uncertainties from 95% confidence bounds. This pressure-derived value of $Q_{c}$ agrees with our image-derived estimate within the experimental uncertainty.